The base of the triangle is 5 cm greater than the height. The area is 33 cm^2. Find the height and the length of the base.
The height of the triangle is ___ cm
The base of the triangle is ___ cm

Answer by swincher4391(1107)

A = 1/2 * B * H
33 = 1/2 * B * H
B = H + 5
33 = (1/2)* (H+5) * H (don't like fractions, multiply by 2)
66 = (H+5)*H
66 = H^2 + 5H
H^2 +5H - 66 = 0
(H+11)(H-6) = 0
H = -11 H = 6 <- we can't have a negative height, so h = 6
Then B = 6 + 5 = 11

The height is 6 cm.
The base is 11 cm.